THE PURE ROTATIONAL SPECTRUM OF CHLORINE NITRATE

\underline{REBECCA A.H. BUTLER}, SIEGHARD ALBERT, Department of Physics, The Ohio State University, 174 West 18th Avenue, Columbus, OH 43210; PAUL HELMINGER, Department of Physics and Chemistry, University of South Alabama, Mobile, Alabama 36688; FRANK C. DE LUCIA, Department of Physics, The Ohio State University, 174 West 18th Avenue, Columbus, OH 43210.

Chlorine nitrate is a molecule of intermediate size whose infrared spectrum is partially resolvable in the Doppler limit. Its Doppler-limited pure rotational spectrum in the microwave is correspondingly dense, not only because of its relatively small rotational constants, but also because it has a number of relatively low-lying vibrational modes. We have recorded this complex spectrum in the microwave spectral region with a FASSST spectrometer. We will report analysis of both ground and excited vibrational states in both the ^{35}Cl and ^{37}Cl species.